# New Insights Into Achievable Spectral Efficiency With Adaptive Free Space Optical Transmissions

### v8

@@ III-C Low-SNR and High-SNR Asymptotic Expressions
    • A common inconsistency in the literature is the omission of the parameter $A_{0}$ when modeling optical fading without pointing errors. While $A_{0}=1$ is often assumed for simplicity, practical FSO systems may exhibit $A_{0}<1$ due to beam divergence and finite aperture effects. Both cases are realistic in practice, but neglecting $A_{0}<1$, when not physically justified, can lead to an overestimated performance, as illustrated in Fig. 4.
      From Fig. 4, the degradation in ASE due to varying turbulence (weak to strong) under the $A_{0}=1$ condition is modest at only $0.5$ bits/s/Hz. In contrast, when $A_{0}<1$, the ASE degradation due to turbulence increases significantly to 2.6 bits/s/Hz. This contrast highlights the importance of modeling $A_{0}$ accurately. The vertical downward-pointing arrows in Fig. 4 further indicate that the ASE reduction associated with $A_{0}<1$ is more severe under strong turbulence and less pronounced under weak turbulence.
@@
      where $w_{\mathrm{L_{eq}}}$ denotes the received equivalent beam waist (see just above Eq. (5)). As turbulence intensity increases, $w_{\mathrm{L_{eq}}}$ also increases, and vice versa. Hence, the ASE loss associated with a given increment in pointing jitter exhibits an inverse dependence on turbulence strength: the penalty is reduced under stronger turbulence fluctuations and amplified under weaker fluctuations, as illustrated by the vertical down-arrowed lines in Fig. 2.
    Figure 5: Spectral efficiency at low SNRs for the GG turbulence channel without pointing jitter under HD detection scheme (see Table IV for turbulence settings).
@@
    Discussion: Fig. 5 presents the exact ASE for the GG fading channel without pointing error under the HD detection scheme at low SNRs. Somewhat surprisingly, the low-SNR spectral efficiency improves with increasing turbulence. This agrees with (86), which shows that strengthening GG turbulence from profile $(a_{i},b_{i})$ to $(a_{j},b_{j})$ yields the following ASE improvement factor at sufficiently low SNRs:
    $\text{ASE improvement factor}\,\approx\,\frac{a_{j}b_{j}}{a_{i}b_{i}}.$ (127)
    This ASE improvement with turbulence at low SNRs is attributed to the enhanced distribution of higher channel gains with increasing turbulence, as shown in Fig. 6. At low SNR, the transmitter, assuming full CSI, adapts the beam power optimally, allocating more power to stronger channel gains while avoiding power wastage on weaker states. Consequently, power adaptation yields a net ASE gain at low SNRs, as confirmed by the numerical results in Fig. 5. For instance, at a target SNR of 10-10 dB/Hz, the ASE improves from $0.1816$ bits/s/Hz under weak turbulence to $0.2434$ bits/s/Hz under strong turbulence ($34\%$ increase). In a wavelength-division multiplexed (WDM) system with $50$ GHz optical bandwidth per channel, this corresponds to an increase from about $9.08$ Gb/s to $12.172$ Gb/s per channel for the same transmit power budget. Such per-channel low-SNR conditions are not typical in current deployments but may plausibly arise in future ultra-wideband WDM coherent FSO systems where a fixed eye-safety-limited transmit power must be shared among many carriers, with atmospheric variability (e.g., haze or fog) further accentuating this effect. At a lower target SNR of 15-15 dB/Hz, the ASE increases from about $3.52$ Gbps/channel to $5.42$ Gbps/channel ($53\%$ gain). Moreover, the channel’s ASE per unit SNR-known as the wideband slope [51]-improves as the SNR decreases: under strong turbulence it rises from $2.43$ units at 10-10 dB to $3.43$ units at 15-15 dB ($41\%$ increase), while under weak turbulence it increases from $1.82$ to $2.27$ units ($25\%$ gain). [^1] Here, ‘unit’ denotes bits/s/Hz per unit SNR, with SNR expressed in linear ratio.
@@
    Figure 6: The PDF of the GG fading gain for the settings described in Table IV. Notice the larger probabilities for the higher gains with increasing turbulence.
@@
    Figure 7: Spectral efficiency results in the low SNR regime for gamma-gamma turbulence channel with ‘mild’ pointing jitter under HD detection scheme.
@@
    In contrast, the effect of turbulence on the spectral efficiency of the optical channel with pointing error at low SNRs is more involved. As depicted in Figures 8-7, the spectral efficiency initially shows a slight improvement in the weak-to-moderate turbulence regime, but then begins to degrade as turbulence strength increases from moderate to strong levels.
@@
    Figure 8: Spectral efficiency results in the low SNR regime for gamma-gamma turbulence channel with ‘strong’ pointing jitter under HD detection scheme.
@@
    This unusual behavior can be explained using the asymptotic low-SNR spectral efficiency result in (78) from Theorem 2. The combined effect of atmospheric turbulence and pointing error on the optical channel’s spectral efficiency at low SNR is encapsulated in the term $A_{0}/4ab$, which can be rewritten as
@@
    $\textrm{Scaling factor}\,\triangleq\,A_{0}\times\dfrac{1}{4ab}.$ (128)
    Figure 9: Variation of the turbulence scaled factor for the GG turbulence channel under HD detection scheme.
@@
    Fig. 9 illustrates the variation of the scaling factor in (128) (red curve) with turbulence, alongside the factor 1/4ab{1}/{4ab} (blue curve), which applies to pure optical turbulence (see (86)). A comparison of the two curves reveals that the parameter $A_{0}$, representing the fraction of collected power, decreases with turbulence-initially slowly in the weak fluctuation regime and then more rapidly under strong turbulence, while the factor 1/4ab{1}/{4ab} increases. For further insight into how $A_{0}$ varies with turbulence, refer to the $A_{0}$ values in Table IV, which provide typical values for weak, moderate, and strong turbulence conditions. In general, except under weak fluctuations, the spectral efficiency of the channel at low SNR with both fading and pointing error tends to degrade with increasing turbulence over a broad range of moderate-to-strong conditions.
@@
    Interestingly, the asymptotic result (78) indicates that the influence of pointing jitter strength $\xi^{2}$ on the channel’s ASE vanishes in the low-SNR regime: the transmitter operates only during channel peaks, and it can be verified that the distribution of these peaks is largely insensitive to $\xi^{2}$. The low-SNR ASE characterization for mobile RF channels in [52] reveals two key insights: power adaptation reduces to an asymptotically optimal on–off strategy, and transmitter feedback collapses to a single bit information. These simplifications are expected to carry over to terrestrial optical channels at low SNRs, offering practical benefits for FSO system design.
@@ V Conclusion
    The proposed exact solutions capture the impact of fading and pointing error parameters, with particular emphasis on the high and low SNR regimes, and highlight an interplay between $A_{0}$ and $\xi^{2}$ as turbulence and pointing error conditions vary. Under mild or negligible pointing error condition-typical of practical systems with well-designed ATP-the expected degradation of spectral efficiency with increasing turbulence is observed. In contrast, for channels with strong pointing error condition-typical of systems with poorly designed or absent ATP, which may occur in low-cost deployments-a notable improvement in spectral efficiency is observed at high SNRs as turbulence increases from weak conditions, since the reduction in pointing jitter outweighs the loss in received power. Overall, these results provide insights across the full spectrum of ATP capabilities, from practical well-designed systems to theoretical or low-cost configurations with no or poorly performing ATP.

### v9

@@ IV Numerical Results and Discussion
  • A common inconsistency in the literature is the omission of the parameter $A_{\mathrm{\scriptscriptstyle 0}}$ when modeling optical fading without pointing errors. While $A_{\mathrm{\scriptscriptstyle 0}}=1$ is often assumed for simplicity, practical FSO systems may exhibit $A_{\mathrm{\scriptscriptstyle 0}}<1$ due to beam divergence and finite aperture effects. Both cases are realistic in practice, but neglecting $A_{\mathrm{\scriptscriptstyle 0}}<1$, when not physically justified, can lead to an overestimated performance, as illustrated in Fig. 4.
    From Fig. 4, the degradation in ASE due to varying turbulence (weak to strong) under the $A_{\mathrm{\scriptscriptstyle 0}}=1$ condition is modest at only $0.5$ bits/s/Hz. In contrast, when $A_{\mathrm{\scriptscriptstyle 0}}<1$, the ASE degradation due to turbulence increases significantly to 2.6 bits/s/Hz. This contrast highlights the importance of modeling $A_{\mathrm{\scriptscriptstyle 0}}$ accurately. The vertical downward-pointing arrows in Fig. 4 further indicate that the ASE reduction associated with $A_{\mathrm{\scriptscriptstyle 0}}<1$ is more severe under strong turbulence and less pronounced under weak turbulence.
@@
    where $w_{{\mathrm{\scriptscriptstyle L_{eq}}}}$ denotes the received equivalent beam waist (see just above Eq. (5)). As turbulence intensity increases, $w_{\mathrm{L_{eq}}}$ also increases, and vice versa. Hence, the ASE loss associated with a given increment in pointing jitter exhibits an inverse dependence on turbulence strength: the penalty is reduced under stronger turbulence fluctuations and amplified under weaker fluctuations, as illustrated by the vertical down-arrowed lines in Fig. 2.
Figure 5: Spectral efficiency at low SNRs for the GG turbulence channel without pointing jitter under the HD detection scheme (see Table I for turbulence settings).
@@
Discussion: Fig. 5 presents the exact ASE for the GG fading channel without pointing error under the HD detection scheme at low SNRs. Somewhat surprisingly, the low-SNR spectral efficiency improves with increasing turbulence. This agrees with (97), which shows that strengthening GG turbulence from profile $(a_{i},b_{i})$ to $(a_{j},b_{j})$ yields the ASE improvement at sufficiently low SNRs, determined by the ratio of the corresponding pre-factors, which can be expressed as
$\text{Relative pre-factor}\,\approx\,\frac{a_{j}b_{j}}{a_{i}b_{i}}.$ (104)
@@
Figure 6: The PDF of the GG fading gain for the settings described in Table I. Notice the larger probabilities for the higher gains with increasing turbulence strength.
@@
Figure 7: Average spectral efficiency results in the low SNR regime for the GG turbulence channel with mild pointing jitter under the coherent HD detection scheme.
This ASE improvement with turbulence at low SNRs is attributed to the enhanced distribution of higher channel gains with increasing turbulence, as shown in Fig. 6. At low SNR, the transmitter, assuming full CSI, adapts the beam power optimally, allocating more power to stronger channel gains while avoiding power wastage on weaker states. Consequently, power adaptation yields a net ASE gain at low SNRs, as confirmed by the numerical results in Fig. 5. For instance, at a target SNR of 10-10 dB/Hz, the ASE improves from $0.1816$ bits/s/Hz under weak turbulence to $0.2434$ bits/s/Hz under strong turbulence ($34\%$ increase). In a wavelength-division multiplexed (WDM) system with $50$ GHz optical bandwidth per channel, this corresponds to an increase from about $9.08$ Gbps to $12.172$ Gbps per channel for the same transmit power budget. Such per-channel low-SNR conditions are not typical in current deployments but may plausibly arise in future ultra-wideband WDM coherent FSO systems where a fixed eye-safety-limited transmit power must be shared among many carriers, with atmospheric variability (e.g., haze or fog) further accentuating this effect. At a lower target SNR of 15-15 dB/Hz, the ASE increases from about $3.52$ Gbps/channel to $5.42$ Gbps/channel ($53\%$ gain). Moreover, the channel’s ASE per unit SNR-known as the wideband slope [49]-improves as the SNR decreases: under strong turbulence it rises from $2.43$ units at 10-10 dB to $3.43$ units at 15-15 dB ($41\%$ increase), while under weak turbulence it increases from $1.82$ to $2.27$ units ($25\%$ gain). Here, ‘unit’ denotes bits/s/Hz per unit SNR, with SNR expressed in linear ratio.
@@
In contrast, the effect of turbulence on the spectral efficiency of the optical channel with pointing error at low SNRs is more involved. As depicted in Figures 7-8, the spectral efficiency initially shows a slight improvement in the weak-to-moderate turbulence regime, but then begins to degrade as turbulence strength increases from moderate to strong levels.
@@
Figure 8: Average spectral efficiency results in the low SNR regime for the GG turbulence channel with strong pointing jitter under the coherent HD detection scheme.
@@
This unusual behavior can be explained using the asymptotic low-SNR spectral efficiency result in (90) from Theorem 2. The net impact of atmospheric turbulence and pointing error on the optical channel’s ASE at low SNR is captured by the scaling coefficient $A_{\mathrm{\scriptscriptstyle 0}}/4ab$, which we rewrite as
@@
$\textrm{ASE scaling factor}\,\triangleq\,A_{\mathrm{\scriptscriptstyle 0}}\times\dfrac{1}{4ab}.$ (105)
@@
Fig. 9 illustrates the variation of the pre-factor in (105) (red curve) valid for turbulence with pointing error, alongside the factor 1/4ab{1}/{4ab} (blue curve), which applies to pure optical turbulence (see (97)). A comparison of the two curves reveals that the parameter $A_{\mathrm{\scriptscriptstyle 0}}$, representing the fraction of collected power, decreases with turbulence-initially slowly in the weak fluctuation regime and then more rapidly under strong turbulence, while the factor 1/4ab{1}/{4ab} increases. For further insight into how $A_{\mathrm{\scriptscriptstyle 0}}$ varies with turbulence, refer to the $A_{\mathrm{\scriptscriptstyle 0}}$ values in Table I, which provide typical values for weak, moderate, and strong turbulence conditions. In general, except under weak fluctuations, the spectral efficiency of the channel at low SNR with both fading and pointing error tends to degrade with increasing turbulence over a broad range of moderate-to-strong conditions.
Figure 9: Variation of the low-SNR ASE pre-factor for the GG turbulence channel with and without pointing jitters under the coherent HD detection scheme.
@@
Interestingly, the asymptotic result (90) indicates that the influence of pointing jitter strength $\xi^{2}$ on the channel’s ASE vanishes in the low-SNR regime: the transmitter operates only during channel peaks, and it can be verified that the distribution of these peaks is largely insensitive to $\xi^{2}$. The low-SNR ASE characterization for mobile RF channels in [50] reveals two key insights: power adaptation reduces to an asymptotically optimal on–off strategy, and transmitter feedback collapses to a single bit information. These simplifications are expected to carry over to terrestrial coherent optical channels at low SNRs, offering practical benefits for FSO system design.
@@ V Conclusion
The proposed exact solutions capture the impact of fading and pointing error parameters, with particular emphasis on the high and low SNR regimes, and highlight an interplay between $A_{\mathrm{\scriptscriptstyle 0}}$ and $\xi^{2}$ as turbulence and pointing error conditions vary. Under mild or negligible pointing error condition-typical of practical systems with well-designed ATP-the expected degradation of spectral efficiency with increasing turbulence is observed. In contrast, for channels with strong pointing error condition-typical of systems with poorly designed or absent ATP, which may occur in low-cost deployments-a notable improvement in spectral efficiency is observed at high SNRs as turbulence increases from weak conditions, since the reduction in pointing jitter outweighs the loss in received power. Overall, these results provide insights across the full spectrum of ATP capabilities, from practical well-designed systems to theoretical or low-cost configurations with no or poorly performing ATP.